OpenStudy (anonymous):

It costs $1 for a plain pretzel and the expense is .60, but customers want more salt which will cost an extra $.25 to make per pretzel. In the past, 100 pretzels were sold calculate how many additional pretzels need to be sold to cover the costs of adding the extra expense of the salt.

OpenStudy (anonymous):

266.66 more pretzels is the answer i think 100 pretzels gives a profit of 40$ Therefore the formula is x= the number of pretzels sold x-(x)expense = profit 100-100(.60)= 40$ then to make 40$ profit, x-(x).85= 40 .15x=40 x=40/.15 and finally it would take 266.7 more pretzels to = 40$ revenue again Am I right?

OpenStudy (dumbcow):

how much do they sell each pretzel for? $1.40

OpenStudy (anonymous):

1$ is the price, .60 is the initial expense sorry messed the question

OpenStudy (dumbcow):

ok then, you are correct that it will take selling 266.66 or 267 pretzels to get profit of $40 but that is only 167 more than original 100

OpenStudy (anonymous):

thanks so much, i hate it when there is no answer key lol. I had to confrim or i wouldnt be able to sleep

OpenStudy (dumbcow):

no problem